OpenStudy (anonymous):

help with another question please??? i don't know what's wrong with me today but i'm just not getting anything -.-

OpenStudy (anonymous):

For which function is the instantaneous rate of change closest to 5 at x=2? a) f(x)=x² b) g(x)=x²+1 c) h(x)=x²+3x-4 d) j(x)=1.1^x

OpenStudy (amistre64):

do you know what a derivative is?

OpenStudy (amistre64):

a and b are equal ... so unless you can pick more than one option they are out

OpenStudy (amistre64):

2x+3 and 1.1^x ln(1.1) are the other 2